OpenStudy (anonymous):

Find the discriminant & describe the type of roots 6x^2-2x-4=0 Show Work!

OpenStudy (anonymous):

@Kashan

OpenStudy (anonymous):

Disc= b^2 - 4ac = (-2)^2 - 4(6)(-4) = 4+64 = 68 Disc>0 so real and distinct roots

OpenStudy (anonymous):

@annawilliams40285 got it..??

OpenStudy (anonymous):

I think it is 96 not 68. But still the nature of roots is correct.. :)

OpenStudy (anonymous):

O sorry.. mistake in calculation.. but result is correct... :)

OpenStudy (anonymous):

@ kashan Can you show me the correct steps to get 96

OpenStudy (anonymous):

Disc= b^2 - 4ac = (-2)^2 - 4(6)(-4) = 4+96 = 100 Disc>0

OpenStudy (anonymous):

@annawilliams40285 got it now..??